# SuperDiff: Diffusion Models for Conditional Generation of Hypothetical New Families of Superconductors

### v2

@@ III.2 Formation Energy
SuperDiff Version Raw % Superconduting True % Superconduting Mean $T_{c}$ Standard Dev. Max $T_{c}$
Entirety of SuperCon 67.81% 62.22% $8.51\,\mathrm{K}$ $8.79\,\mathrm{K}$ $97.0\,\mathrm{K}$
Cuprates 64.53% 58.22% $64.68\,\mathrm{K}$ $14.61\,\mathrm{K}$ $110.5\,\mathrm{K}$
Pnictides 95.74% 96.39% $22.00\,\mathrm{K}$ $2.43\,\mathrm{K}$ $29.5\,\mathrm{K}$
Others 55.82% 47.56% $6.33\,\mathrm{K}$ $2.50\,\mathrm{K}$ $30.5\,\mathrm{K}$
Table 3: The percentages of generated compounds determined to be superconducting by the classification model for each of the unconditional SuperDiff versions, along with the estimated true percentages according to Eq. 7 and summary $T_{c}$ statistics from the predictions by the regression model.
@@
As shown in the figure, unconditional SuperDiff generated a majority of compounds with negative formation energy for all classes of superconductors, with the mean formation energy for all classes predicted to be negative as well. In Ref. [28], it was stated that negative formation energy values for compounds are a good indicator of their stability and synthesizability; therefore, although these predictions are not definitive proof—experimentation validation would be necessary—these predictions provide an indication that most of the compounds generated by unconditional SuperDiff are plausibly stable and synthesizable.
Figure 5: Clustering of the valid generated compounds from the Cuprates version of unconditional SuperDiff. Full circles represent Cuprate data points from SuperCon, and black open circles are generated compounds. We notice that unconditional SuperDiff did not generate any new families of superconductors, as all generated compounds fall within the existing clusters of superconductors from SuperCon.
Figure 6: Clustering of the valid generated compounds from the Others version of conditional SuperDiff conditioned on various different compounds. Full circles represent Others data points from SuperCon, and black open circles are generated compounds. Although some generated superconductors fall within the existing SuperCon clusters, we were able to identify two new clusters consisting of only generated superconductors (marked with arrows). These two new clusters correspond to two new families of superconductors generated by SuperDiff: $\mathrm{Li_{1-x}Be_{x}Ga_{2}Rh}$ and $\mathrm{Na_{1-x}Al_{1-y}Mg_{x+y}Ge_{1-z}Ga_{z}}$.
For conditional SuperDiff, the distribution of formation energies for generated compounds is heavily dependent on the reference. However, given a reasonable reference compound—that is, a valid reference compound that belongs to the class of superconductor that the version of SuperDiff was trained on—we demonstrate that conditional SuperDiff is able to generate compounds predicted to be stable by ElemNet. Specifically, as shown in Fig. 4, for the cuprates version of conditional SuperDiff conditioned on $\mathrm{YBa_{1.4}Sr_{0.6}Cu_{3}O_{6}Se_{0.51}}$ [24] and the pnictides version of conditional SuperDiff conditioned on $\mathrm{BaFe_{1.7}Ni_{0.3}As_{2}}$ [32]—some of the compounds we conditioned conditional SuperDiff on to find new families of superconductors in Section III.5—the predicted distribution of formation energies for generated compounds show all generated compounds to have negative formation energy. These results indicate that, given reasonable reference compounds, conditional SuperDiff can generate plausibly stable and synthesizable compounds, which is not surprising given the fundamental architecture similarities between conditional and unconditional SuperDiff.
@@ III.3 Superconductivity
After those general checks, we performed some computational checks for superconductivity in order to verify that unconditional SuperDiff is indeed able to generate probable superconductors. We ran the compounds generated by unconditional SuperDiff through the $K$-Nearest Neighbors (KNN) classification model and regression model from Ref. [4] for predicting superconductivity and critical temperature, respectively, based on elemental composition.
@@
where $\textit{{tp}}=98.69\%$ and $\textit{{fp}}=16.94\%$ are reported by Ref. [4].
For the generated compounds that were predicted to be superconducting, we used the regression model in Ref. [4] to predict their critical temperatures. Like all other tests done so far, this computational prediction is only an approximation. We tabulated the results of the classification and regression predictions on the compounds generated by unconditional SuperDiff in Table 3. We will discuss the predicted superconductivity of compounds generated by conditional SuperDiff later in Section III.5.
As seen in the table, all versions of unconditional SuperDiff were able to generate predicted superconductors at a rate comparable to the GAN in Ref. [6] and much higher than the 3% achieved by manual search in Ref. [2]—notably, unconditional SuperDiff seems to perform much better on pnictides despite the small training set. This is further indication of the effectiveness of computational search for superconductors when compared to manual searches. Moreover, unconditional SuperDiff seems to capture the critical temperature distribution of the SuperCon training dataset much better than the GAN in Ref. [6].
Although actual synthesis and testing in a lab are required to confirm superconductivity, these checks, combined with the clustering analysis results that we will discuss in Section III.4, provide a general indication that unconditional SuperDiff is able to generate highly plausible superconductors.
@@ III.4 Clustering Results
We ran the clustering analysis described in Section II.5 on both unconditional and conditional SuperDiff. We display the clustering results for the cuprates version of unconditional SuperDiff in Fig. 5. Superconductors from the SuperCon database are shown with full circles of different colors, whereas our predictions are shown with open black circles. Although unconditional SuperDiff generated compounds in all known clusters or families of superconductors, no new families of superconductors were generated by unconditional SuperDiff—this was true for the other versions of unconditional SuperDiff as well. This was the expected result for unconditional SuperDiff as the underlying DDPM’s goal is to just find a mapping from Gaussian noise to the training data distribution, not some other new distribution. However, superconductor discovery has a particular interest in the generation of new families of superconductors, so a method to control the generation process to change the generated data distribution is desirable. With conditional SuperDiff, we are able to control the generation process to computationally generate new families of superconductors for the first time.
In Fig. 6, we display a sample clustering result from the “others” version of conditional SuperDiff conditioned on various compounds. As seen in the plot, we identified two new clusters: $\mathrm{Li_{1-x}Be_{x}Ga_{2}Rh}$, which was generated by conditioning SuperDiff on $\mathrm{LiGa_{2}Rh}$ [36], and $\mathrm{Na_{1-x}Al_{1-y}Mg_{x+y}Ge_{1-z}Ga_{z}}$, which was generated by conditioning SuperDiff on $\mathrm{NaAlGe}$ [37]. Those and other predicted families will be discussed in more detail below.
These clustering results show that, with this ability to control generation, and by conditioning SuperDiff on compounds not in the SuperCon training set, SuperDiff is able to use information from various reference compounds to generate completely new families of superconductors. As expected, we note that for these generated new families, the reference compound does belong to the new cluster; however, the contribution presented in this work is that we are able to extrapolate a new family of superconductors from the single reference compound. We performed this clustering analysis on all versions of conditional SuperDiff conditioned on a variety of different reference compounds, and we discuss the promising new families of superconductors generated by conditional SuperDiff in more detail and verify their superconductivity in Section III.5.
@@ III.5 Promising Generated New Families
Figure 7: Plot of $T_{c}$ predicted by the regression model in Ref. [4] versus Cesium content ($x$) for $\mathrm{Ba_{2-x}Cs_{x}CuO_{3.3}}$ family generated by conditional SuperDiff (see Table 4). We notice a characteristic parabolic dependence of $T_{c}$ versus doping, observed previously in other cuprate families [38].
After running clustering analysis for the different versions of conditional SuperDiff conditioned on a variety of reference compounds, we manually identified the most promising new families of superconductors generated by conditional SuperDiff. Beyond the novelty, uniqueness, and SMACT checks, we further checked for the novelty of these newly generated families by searching on the internet and through other databases—these newly generated families could not be found anywhere else. We tabulated these most promising new families of superconductors generated by conditional SuperDiff in Table 4. There, we identified the reference compound used as well as a few output examples and their respective predicted $T_{c}$ using the regression model in Ref. [4], and determined the general formula for the new family. We notice that most compounds generated with conditional SuperDiff are predicted to be superconducting, with predicted $T_{c}$ being reasonable for each class. Additionally, a particularly interesting result to note was that our model seemed to generate some new families of superconductors with double or, in one case, even triple doping. This is an interesting new avenue for superconductor discovery that has not been extensively studied, which our model suggests should be explored in more detail by material scientists.
We further demonstrate that conditional SuperDiff is able to generate realistic new families of superconductors by plotting the predicted $T_{c}$ using the regression model in Ref. [4] against the Cesium doping content for the newly generated $\mathrm{Ba_{2-x}Cs_{x}CuO_{3.3}}$ family in Fig. 7. We notice that the generated $\mathrm{Ba_{2-x}Cs_{x}CuO_{3.3}}$ family is predicted to exhibit the expected parabolic $T_{c}$ doping dependence relationship for this type of cuprate superconductor, which was observed previously in other cuprate families [38].
@@ IV Conclusion
In this paper, we have introduced a novel method for superconductor discovery using diffusion models with conditioning functionality that has addressed these major issues. Like previous works applying generative models to superconductor discovery, we were able to generate novel, realistic, and highly plausible superconductors that lie outside of existing databases—leveraging this “inverse design” approach to significantly outperform manual search and previous classification model approaches. With our unconditional model, we were also able to address the low generated compound uniqueness issues that plagued previous works due to the small training data set for pnictides. Most importantly, however, beyond the unconditional performance improvements the diffusion model brought, our contribution of implementing conditioning with ILVR for superconductor discovery to allow the generation process to be controlled enabled the creation of a tool for computationally generating completely new families of superconductors. We verified the generation of new families of superconductors with our clustering analysis, and we presented several of these promising new families of generated superconductors for several different classes of superconductors in Table 4. Once again, we point out that no previous computational model for superconductor discovery would have been capable of generating these new families of superconductors, as they attempt to produce only samples that match the training data.
@@
The application of deep generative models for superconductor discovery continues to be a very promising and exciting approach. Future studies can benefit from possible improvements that can be made to SuperDiff, including implementing a physics-informed diffusion model and creating and utilizing a better, more comprehensive training dataset of superconductors. Nevertheless, SuperDiff in its current form is still very powerful as a tool for superconductor discovery, and researchers can currently benefit from it in a myriad of ways, such as by using its novel generations as inspiration—starting with the new families introduced here, using it to expand on their own new discoveries, or by simply experimenting with many more reference compounds to continue using it to generate completely new families of superconductors.

### v3

@@ III.2 Formation Energy
As shown in the figure, unconditional SuperDiff generated a majority of compounds with negative formation energy for all classes of superconductors, with the mean formation energy for all classes predicted to be negative as well. In Jha et al. [28], it was stated that negative formation energy values for compounds are a good indicator of their stability and synthesizability; therefore, although these predictions are not definitive proof—experimentation validation would be necessary—these predictions provide an indication that most of the compounds generated by unconditional SuperDiff are plausibly stable and synthesizable.
Figure 5: Clustering of the (a) valid generated compounds from the Cuprates version of unconditional SuperDiff and (b) valid generated compounds from the Others version of conditional SuperDiff conditioned on various different compounds. Colored full circles represent data points from SuperCon (cuprates only for (a) and others only for (b)), with each color representing a different cluster, or family, of superconductor as identified by the model from Roter et al. [15]; black open circles are compounds generated by SuperDiff. We notice that unconditional SuperDiff did not generate any new families of superconductors, as all generated compounds fall within the existing clusters of superconductors from SuperCon. However, for conditional SuperDiff, although some generated superconductors fall within the existing SuperCon clusters, we were able to identify two new clusters consisting of only generated superconductors (marked with arrows). These two new clusters correspond to two new families of superconductors generated by SuperDiff: $\mathrm{Li_{1-x}Be_{x}Ga_{2}Rh}$ and $\mathrm{Na_{1-x}Al_{1-y}Mg_{x+y}Ge_{1-z}Ga_{z}}$.
For conditional SuperDiff, the distribution of formation energies for generated compounds is heavily dependent on the reference compound. However, given a reasonable reference compound—that is, a valid reference compound that belongs to the class of superconductor that the version of SuperDiff was trained on—we demonstrate that conditional SuperDiff is able to generate compounds predicted to be stable by ElemNet. Specifically, as shown in Fig. 4, for the cuprates version of conditional SuperDiff conditioned on $\mathrm{YBa_{1.4}Sr_{0.6}Cu_{3}O_{6}Se_{0.51}}$ [24] and the pnictides version of conditional SuperDiff conditioned on $\mathrm{BaFe_{1.7}Ni_{0.3}As_{2}}$ [32]—some of the compounds we conditioned conditional SuperDiff on to find new families of superconductors later—the predicted distribution of formation energies for generated compounds show all generated compounds to have negative formation energy. These results indicate that, given reasonable reference compounds, conditional SuperDiff can generate plausibly stable and synthesizable compounds, which is not surprising given the fundamental architecture similarities between conditional and unconditional SuperDiff.
@@ III.3 Superconductivity
After those general checks, we performed some computational checks for superconductivity in order to verify that unconditional SuperDiff is indeed able to generate probable superconductors. We ran the compounds generated by unconditional SuperDiff through the $K$-Nearest Neighbors (KNN) classification model and regression model from Roter and Dordevic [4] for predicting superconductivity and critical temperature, respectively, based on elemental composition.
@@
where $\textit{{tp}}=98.69\%$ and $\textit{{fp}}=16.94\%$ are reported by Roter and Dordevic [4].
For the generated compounds that were predicted to be superconducting, we used the regression model in Roter and Dordevic [4] to predict their critical temperatures. Like all other tests done so far, this computational prediction is only an approximation. We tabulated the results of the classification and regression predictions on the compounds generated by unconditional SuperDiff in Table 1. We will discuss the predicted superconductivity of compounds generated by conditional SuperDiff later.
As seen in the table, all versions of unconditional SuperDiff were able to generate predicted superconductors at a rate comparable to the GAN in Kim and Dordevic [6] and much higher than the 3% achieved by manual search in Hosono et al. [2]—notably, unconditional SuperDiff seems to perform much better on pnictides despite the small training set. This is further indication of the effectiveness of computational search for superconductors when compared to manual searches. Moreover, unconditional SuperDiff seems to capture the critical temperature distribution of the SuperCon training dataset much better than the GAN in Kim and Dordevic [6].
Although actual synthesis and testing in a lab are required to confirm superconductivity, these checks, combined with the clustering analysis results that we will discuss later, provide a general indication that unconditional SuperDiff is able to generate highly plausible superconductors.
@@ III.4 Clustering Results
We ran the clustering analysis described previously on both unconditional and conditional SuperDiff. We display the clustering results for the cuprates version of unconditional SuperDiff in Fig. 5. Superconductors from the SuperCon database are shown with full circles of different colors, whereas our predictions are shown with open black circles. Although unconditional SuperDiff generated compounds in all known clusters or families of superconductors, no new families of superconductors were generated by unconditional SuperDiff—this was true for the other versions of unconditional SuperDiff as well. This was the expected result for unconditional SuperDiff as the underlying DDPM’s goal is to just find a mapping from Gaussian noise to the training data distribution, not some other new distribution. However, superconductor discovery has a particular interest in the generation of new families of superconductors, so a method to control the generation process to change the generated data distribution is desirable. With conditional SuperDiff, we are able to control the generation process to computationally generate new families of superconductors for the first time.
In Fig. 5, we also display a sample clustering result from the “others” version of conditional SuperDiff conditioned on various compounds. As seen in the plot, we identified two new clusters: $\mathrm{Li_{1-x}Be_{x}Ga_{2}Rh}$, which was generated by conditioning SuperDiff on $\mathrm{LiGa_{2}Rh}$ [36], and $\mathrm{Na_{1-x}Al_{1-y}Mg_{x+y}Ge_{1-z}Ga_{z}}$, which was generated by conditioning SuperDiff on $\mathrm{NaAlGe}$ [37]. Those and other predicted families will be discussed in more detail below.
These clustering results show that, with this ability to control generation, and by conditioning SuperDiff on compounds not in the SuperCon training set, SuperDiff is able to use information from various reference compounds to generate completely new families of superconductors. As expected, due to the nature of the conditioning method, we note that for these generated new families, the reference compound does belong to the new cluster generated based on it; however, one of the main contributions presented in this work is that we are able to extrapolate a new family of superconductors from an otherwise single reference compound. We performed this clustering analysis on all versions of conditional SuperDiff conditioned on a variety of different reference compounds, and we discuss the promising new families of superconductors generated by conditional SuperDiff in more detail and verify their superconductivity below.
@@ III.5 Promising Generated New Families
Figure 6: Plot of $T_{c}$ predicted by the regression model in Roter and Dordevic [4] versus Cesium content ($x$) for $\mathrm{Ba_{2-x}Cs_{x}CuO_{3.3}}$ family generated by conditional SuperDiff (see Table 2). We notice a characteristic parabolic dependence of $T_{c}$ versus doping, observed previously in other cuprate families [38].
After running clustering analysis for the different versions of conditional SuperDiff conditioned on a variety of reference compounds, we manually identified the most promising new families of superconductors generated by conditional SuperDiff. Beyond the novelty, uniqueness, and SMACT checks, we further checked for the novelty of these newly generated families by searching on the internet and through other databases—these newly generated families could not be found anywhere else. We tabulated these most promising new families of superconductors generated by conditional SuperDiff in Table 2. There, we identified the reference compound used as well as a few output examples and their respective predicted $T_{c}$ using the regression model in Roter and Dordevic [4], and determined the general formula for the new family. We notice that most compounds generated with conditional SuperDiff are predicted to be superconducting, with predicted $T_{c}$ being reasonable for each class. Additionally, a particularly interesting result to note was that our model seemed to generate some new families of superconductors with double or, in one case, even triple doping. This is an interesting new avenue for superconductor discovery that has not been extensively studied, which our model suggests should be explored in more detail by material scientists.
We further demonstrate that conditional SuperDiff is able to generate realistic new families of superconductors by plotting the predicted $T_{c}$ using the regression model in Roter and Dordevic [4] against the Cesium doping content for the newly generated $\mathrm{Ba_{2-x}Cs_{x}CuO_{3.3}}$ family in Fig. 6. We notice that the generated $\mathrm{Ba_{2-x}Cs_{x}CuO_{3.3}}$ family is predicted to exhibit the expected parabolic $T_{c}$ doping dependence relationship for this type of cuprate superconductor, which was observed previously in other cuprate families [38].
@@ IV Discussion
In this paper, we have introduced a novel method for superconductor discovery using diffusion models with conditioning functionality that has addressed these major issues. Like previous works applying generative models to superconductor discovery, we were able to generate novel, realistic, and highly plausible superconductors that lie outside of existing databases—leveraging this “inverse design” approach to significantly outperform manual search and previous classification model approaches. With our unconditional model, we were also able to address the low generated compound uniqueness issues that plagued previous works due to the small training data set for pnictides. Most importantly, however, beyond the unconditional performance improvements the diffusion model brought, our contribution of implementing conditioning with ILVR for superconductor discovery to allow the generation process to be controlled enabled the creation of a tool for computationally generating completely new families of superconductors. We verified the generation of new families of superconductors with our clustering analysis, and we presented several of these promising new families of generated superconductors for several different classes of superconductors in Table 2. Once again, we point out that no previous computational model for superconductor discovery would have been capable of generating these new families of superconductors as they attempt to produce only samples that match the training data.
@@
The application of deep generative models for superconductor discovery continues to be a very promising and exciting approach. Future studies can benefit from possible improvements that can be made to SuperDiff, including implementing a physics-informed diffusion model and creating and utilizing a better, more comprehensive training dataset of superconductors. Nevertheless, SuperDiff in its current form is still very powerful as a tool for superconductor discovery, and researchers can currently benefit from it in a myriad of ways, such as by using its novel generations as inspiration—starting with the new families introduced here, using it to expand on their own new discoveries, or by simply experimenting with many more reference compounds (such as high-pressure superconductors) to continue using it to generate completely new families of hypothetical superconductors or hypothetical superconductors with even higher $T_{c}$.
Data availability
The SuperCon dataset [21] used in this study to train the SuperDiff model is publicly available at https://doi.org/10.48505/nims.3739, and a copy of the processed dataset used is available at https://github.com/sdkyuanpanda/SuperDiff/tree/54f0520a67bf8308fbf437b2b66aa36beee52acd/datasets. The 265,722 valid compounds generated by the four versions of unconditional SuperDiff and the 270 valid compounds generated by conditional SuperDiff conditioned on $\mathrm{YBa_{1.4}Sr_{0.6}Cu_{3}O_{6}Se_{0.51}}$ [24] are available at https://github.com/sdkyuanpanda/SuperDiff/tree/54f0520a67bf8308fbf437b2b66aa36beee52acd/outputs. Other data that support the results of this study are available from the corresponding author upon reasonable request.
Code availability
An implementation of the proposed model, SuperDiff, is publicly available online at https://github.com/sdkyuanpanda/SuperDiff and is citable on Zenodo at https://doi.org/10.5281/zenodo.10699906 [39].
Additional information
Author contributions statement Conceptualization and design: S.Y. Data analysis and interpretation: S.Y. and S.V.D. Drafting of the manuscript: S.Y. Critical revision of the manuscript for important intellectual content: S.Y. and S.V.D. Supervision: S.V.D. All authors read and approved the final manuscript.
Competing interests The authors declare no competing interests.